OpenStudy (anonymous):

Given the following functions, find f(g(9)). f(x)=1/x-10, g(x) 1/x. I'm really confused so please help me. Thank you.

OpenStudy (anonymous):

\[f(g(9))\] first find \(g(9)=\frac{1}{9}\) then find \(f(\frac{1}{9})\)

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

\[f(x)=\frac{1}{x-10}\] \[f(\frac{1}{9})=\frac{1}{\frac{1}{9}-10}\]

OpenStudy (anonymous):

is this the answer: 89/9?